OpenStudy (anonymous):

What is the product of 4the square root of 6. • 6the square root of 2.? Simplify if possible.

OpenStudy (anonymous):

24the square root of 2. 48the square root of 2. 24the square root of 3. 48the square root of 3.

OpenStudy (anonymous):

Help please @thomas_2016

OpenStudy (anonymous):

Is it written with the radical sign?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

does it want like a reduced radical or what?

OpenStudy (anonymous):

idk but what do you think it is?

OpenStudy (anonymous):

Does it give any other directions?

OpenStudy (anonymous):

I'm taking an online test and have 4 more equations so I really need help

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I gonna guess on this one and post one of the others

OpenStudy (anonymous):

Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

it does come out even.

OpenStudy (anonymous):

Do you know this one?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

Sorry my internet is acting very slow

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

It is all good.

OpenStudy (anonymous):

I have two more

OpenStudy (anonymous):

ok post

OpenStudy (anonymous):

Is the test timed?

OpenStudy (anonymous):

na

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

worng thing

OpenStudy (anonymous):

Evaluate fifth root of 8 multiplied by square root of 8 over cube root of 8 to the power of 5 8 to the power of negative 3 over 10 8 to the power of negative 41 over 30 8 to the power of negative 29 over 30 8 to the power of 71 over 30

OpenStudy (anonymous):

do you know it?

OpenStudy (anonymous):

I am working it out

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is the power a fraction?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

k

OpenStudy (anonymous):

If I don't finish my math in a 7 day period I have to go to public school for 8th grade and everything gets taken away from me (xbox, phone, ipod, laptop, TV, music player and my car) so please help with one more when your done

OpenStudy (anonymous):

Im sorry but I can get I good answer.

OpenStudy (anonymous):

guess then

OpenStudy (anonymous):

3rd